OpenStudy (anonymous):

Members of the Ski Club contributed equally to obtain $1800 for a holiday trip. When 6 members found they could not go, their contributions were refunded and each remaining member then had to pay an additional $10 to remain at their goal of $1800. How many went on the trip?

OpenStudy (anonymous):

x^2 - 6x - 1080 = 0 Using the quadratic equation, x = 36 --- this is the number of original members who signed up Therefore, the number who went to the trip = 36 - 6 = 30 members